OpenStudy (anonymous):

The following polynomial represents a profit function for a certain production line, where x is a number of produced units, find a. the zeros and the multiplicity of each b. where the graph crosses or touches the x-axis c. number of turning points d. explain the meaning of the above, if any f(x)=(x+5)(x−3)(x+3)

OpenStudy (jack1):

graph crosses the x axis at y = 0 y= 0 when either x+5 = 0 x-3 = 0 and/or x+3 = 0 (as anything times 0 = 0) so from that you can get your turning points: x+5 = 0, therefore x=-5 x-3 = 0, therefore x=3 x+3 = 0, therefore x=-3 your function crosses the x axis at the points x= -5, -3, and 3

OpenStudy (jack1):

you following ok so far, do you want anything explained a bit more?

OpenStudy (anonymous):

thank you its just so confusing math hates me lol

OpenStudy (jack1):

sók, do you want me to move on to the next part or do you not understand the above yet and want a bit more detail...?

OpenStudy (anonymous):

no go ahead i will have to go over this about 5 times to understand it but i will get it ;)

OpenStudy (jack1):

ok, this is what it looks like:

OpenStudy (jack1):

so as you can see, it's a cubic polynomial, with 2 turning points to work them out mathematically, multiply out the brackets so you're left with the form: f(x) = ax^3 + bx^2 + cx +d in this case: f(x) = (x+5)(x−3)(x+3) f(x) = x^3 + 5x^2 -9x -45

OpenStudy (jack1):

now to find the turning points: they will occur when the gradient of the graph = 0 the equation for gradient is the derivative of the equation of the function so, the equation for your function: f(x) = x^3 + 5x^2 -9x -45 therefore derivative of function: f '(x) = 3x^2 + 10x -9

OpenStudy (jack1):

now you can use the quadratic formula to find the exact position of when 3x^2 + 10x -9 = 0, which is the x co-ordinates of the turning points in this case it's at: x= -4.07 ish and x= 0.737 ish

OpenStudy (jack1):

so as it's a profitability graph, and x = number of units produced... i'd assume that y = profits made...? maybe...? i'm not sure how to do the explanation part, sorry, just the maths and i don't know what a zero or a multiplicity is...? do you have any ideas?

OpenStudy (anonymous):

multiplicity i believe it is what you multiply like if it had^2 in it it would be 2. or if it had ^3 answer would be 3.as for the zero i dont know